# Danny Halawi4, Fred Zhang4, Philip E. Tetlock2,5

### v4

@@ 3.4 Leaderboards
We generate leaderboards, ranking models and humans by average overall score. The main leaderboard highlights the top forecasting submission across all questions and can by sorted by performance on the source type (market or dataset) and by resolution status (resolved or unresolved questions). The leaderboard is updated nightly, given the latest data, market resolution values, and crowd forecasts.
@@ 3.5 Datasets
Every time we run the public survey outlined in Section 4, we provide multiple independent forecasts and rationales on every one of the 200 forecast questions and report the accuracy of the median public forecast. See Section B.2.1 for details.
@@
We provided a random subset of the benchmark questions to 39 expert forecasters with a strong track record of accurate performance on a diverse set of geopolitical questions. In addition to forecasts and rationale, the superforecasters provide pertinent information about their forecasting process, like search terms used and useful URLs consulted. See Section B.2.2.
@@
Similar to the general public dataset, we prompt LLMs to produce forecasts on each of 1,000 forecast questions in the LLM question set. Their rationales are also included in the dataset whenever provided. See Section B.2.3 for details.
@@
In creating the benchmark, we have automated question creation and resolution from all of the sources outlined in Table 1. We provide these as a dataset that can be combined with the forecast datasets mentioned above. See Section B.1 for details on these datasets.
@@ 4 Human forecaster baseline
To compare LLM forecasting performance to human performance, we ran surveys of two different groups: the general public and expert forecasters (superforecasters).
4.1 Public
We recruited 500 human forecasters via Prolific and Facebook advertising to participate as representatives of the general public. These human subjects completed a brief introductory survey to gather demographic information [^6] See Appendix I for an overview of public participant demographics. and evaluate performance on a few forecasting and comprehension tasks. They then completed a one-hour survey containing 20 random questions from the 200-item human question set described in Section 3.2, providing their forecasts and rationales for each question.
The number of responses per question varied to ensure representativeness across subject matters and sources; at least 40 responses were gathered per question, averaging 49 responses per question.
@@ 4.2 Superforecasters
We additionally recruited 39 superforecasters, who have a strong track record of accurate performance on a diverse set of geopolitical questions, to participate in a 9-day forecasting experiment in which participants were prompted to give their individual forecasts for 20 random questions from the same 200-item human question set described above. Roughly halfway through the 9-day experiment, participants were moved into a group forecasting stage in which we allowed them to see one another’s forecasts and rationales and to communicate about each question. They were also given the opportunity to forecast on questions beyond the 20 questions assigned to them individually.
Because of the more limited quantity of the superforecasters, questions generally had fewer responses than in the public survey; a minimum of 3 forecasts were recorded for each question, with an average of 8 responses per question.
See Figure 2 for an example question from the human surveys. All data from human surveys are anonymized per IRB requirements.
Results
We evaluate human performance on the forecasting tasks by comparing predictions to the ground truth on resolved questions and comparing to community aggregates on unresolved questions. We find that the median public survey participant had an overall Brier score of 0.111 on the 200-item subset we drew from the LLM Question Set with 1,000 questions. The median superforecaster participant had an overall Brier score of 0.091 on this same subset of forecasting questions. See Section 5 and Table 2 for a comparison between human performance and LLM performance, as well as Appendix H for a top 50 leaderboard of human and LLM performance.
@@ 5 LLM baseline
In this section, we evaluate the forecasting capabilities of language models and report on the methodology and results.
@@ 5.1 Methodology
We evaluate a suite of instruction-following chat models without any additional fine-tuning (see Section 2 for details on the models). In each baseline, we prompt the model to generate a probabilistic forecast that the question will resolve to “Yes.”
@@
    Our first baseline, prompts the model zero-shot to generate a forecast directly without generating other content, such as intermediate thinking (Figure 4). By prompting the model to output its forecast directly, we assess raw forecasting capability without sensitivity to prompting strategies.
@@
    Since LLMs’ knowledge are not continuously updated, it is important to provide LLMs with up-to-date information relevant to the question (Zou et al., 2022). Our third baseline, scratchpad with news, uses the same scratchpad prompt as above, supplemented with retrieved news articles. The retrieval system is the same as described in Halawi et al. (2024): an LLM generates search queries for a news API, filters articles for relevancy, and summarizes the articles.
@@
    The question sets we provide to LLMs contain what we term freeze values. For market questions these are just the crowd forecast on the market the day the question set was created, as described in Section 3.2. For dataset questions, these are baseline values relevant to the forecasting task. [^7] For example, in a question generated from a Wikipedia page about whether a chess player’s Elo rating will increase by a given date, the freeze value is the chess player’s Elo rating on the question set generation date. An explanation of what the freeze value represents is also provided. Our third baseline is the zero-shot with freeze values. This is simply the zero-shot prompt from Baseline 1 supplemented with the freeze value and an explanation of the freeze value. For examples of the freeze value its explanation, see Table 8 and Table 9.
@@
    The aggregated prediction of forecasters (“crowd” prediction) is a strong benchmark. For instance, Metaculus (2023) shows that an ensemble of all forecasters consistently outperforms using just the top 5, 10, …, 30 best forecasters (based on past scores). In our final baseline, we aggregate the predictions generated by LLMs into an LLM “crowd” prediction.
    To produce the LLM crowd forecast, we use 3 models (GPT-4o, Claude-3.5-Sonnet, and Gemini-1.5-Pro) and 3 prompts crafted by superforecasters (Figure 6, Figure 7, and Figure 8). This results in 9 forecasts per question. We generate 3 LLM crowd baselines using the median, geometric mean, and geometric mean of log odds (Satopää et al., 2014). For details, see Appendix E.
@@ 5.2 Results
In Table 2, we show that superforecasters achieve an overall mean Brier score of $0.092$, significantly outperforming both the general public (Brier $=0.114$, $p<0.001$) and the top LLM performer on the 200-item subset (Claude 3.5 Sonnet: Brier $=0.114$, $p=0.001$). [^8] See statistical note in Appendix G. The top-performing language models all had access to the crowd forecast on market questions (the “freeze values” from Baselines 4, 5, and 6 above). The top-performing model without access to the crowd forecast on market questions was less accurate than models with access to the human forecast with a Brier score of $0.126$. The comparison between humans and LLMs relies on the 200 questions forecasted by humans, which is a random sub-sample of the 1,000 questions in the question set provided to LLMs (excluding combination questions). [^9] Accuracy measures are based on more than 200 forecasts because human and LLM forecasters submitted multiple forecasts on each dataset question, one for each time horizon. The results presented here include forecasts over the 7-day, 30-day, and 90-day time horizons.
@@
Superforecaster median forecast ForecastBench 0.126 0.059 0.092 [0.073, 0.112] 0%
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Scratchpad 0.139 0.089 0.114 [0.092, 0.136] <0.01 31%
Public median forecast ForecastBench 0.156 0.072 0.114 [0.095, 0.133] <0.001 24%
Claude-3-5-Sonnet-20240620 Anthropic News with freeze values Scratchpad 0.144 0.094 0.119 [0.098, 0.141] <0.001 30%
GPT-4 OpenAI Freeze values Zero shot 0.164 0.079 0.121 [0.100, 0.143] <0.001 32%
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Zero shot 0.145 0.099 0.122 [0.095, 0.149] <0.001 32%
GPT-4-Turbo-2024-04-09 OpenAI Freeze values Zero shot 0.170 0.077 0.124 [0.102, 0.145] <0.001 33%
GPT-4o OpenAI News with freeze values Scratchpad 0.168 0.082 0.125 [0.107, 0.144] <0.001 26%
Claude-3-5-Sonnet-20240620 Anthropic Scratchpad 0.139 0.113 0.126 [0.104, 0.148] <0.001 29%
GPT-4o OpenAI Freeze values Scratchpad 0.163 0.093 0.128 [0.108, 0.148] <0.001 27%
Always 0.5 ForecastBench 0.250 0.184 0.217 [0.207, 0.227] <0.001 16%
@@
1. Shows performance on the 200 standard questions provided in the human question set at the 7-, 30-, and 90-day forecast horizons.
2. The full leaderboard is available at www.forecastbench.org. Online results are updated nightly, so may be slightly different than the static version presented here.
3. For resolved questions, predictions are compared against ground truth while for unresolved questions, they are compared to community aggregates.
@@
5. Pct. more accurate than No. 1: The percent of questions where this forecaster had a better overall score than the best forecaster (with rank 1).
6. Pairwise p-value comparing to No. 1 (bootstrapped): The p-value calculated by bootstrapping the differences in overall score between each model and the best forecaster (the group with rank 1) under the null hypothesis that there’s no difference.
As a particular failure mode, we find LLMs are specifically worse at combination questions. Although our human surveys did not explicitly ask for forecasts on combination questions, we bound human performance by assuming independence of the component of each combination question. This underestimates human accuracy because a human forecaster predicting the outcome of a combination question could account for independence between the permuted events. In Table 20, we present this comparison of human and LLM forecasts. We see that LLMs perform poorly on these combination questions, and including them in the benchmark widens the gap between human and LLM performance: superforecasters (Brier score of $0.071$) outperform the general public (Brier score of $0.090$) and the top LLM (Claude-3.5 Sonnet, Brier score of $0.124$) significantly. To benchmark the size of this gap in performance, the $0.053$ Brier score gap in performance between superforecasters and Claude-3.5 Sonnet (the top-performing LLM) is significantly larger than the $0.035$ gap in performance between Claude-3.5 Sonnet and Claude 2.1.
@@
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Scratchpad 0.171 0.062 0.116 [0.111, 0.122] 0%
GPT-4-Turbo-2024-04-09 OpenAI Freeze values Scratchpad 0.171 0.065 0.118 [0.112, 0.123] 0.243 42%
GPT-4o OpenAI Freeze values Scratchpad 0.191 0.054 0.123 [0.117, 0.128] <0.001 43%
Gemini-1.5-Pro Google Freeze values Scratchpad 0.162 0.089 0.125 [0.120, 0.131] <0.001 35%
GPT-4o OpenAI News with freeze values Scratchpad 0.192 0.066 0.129 [0.123, 0.134] <0.001 39%
Gemini-1.5-Pro Google News with freeze values Scratchpad 0.167 0.094 0.131 [0.125, 0.136] <0.001 34%
Claude-3-Opus-20240229 Anthropic Freeze values Zero shot 0.190 0.076 0.133 [0.126, 0.139] <0.001 41%
Qwen1.5-110B-Chat Qwen Freeze values Scratchpad 0.176 0.092 0.134 [0.128, 0.139] <0.001 31%
GPT-4-Turbo-2024-04-09 OpenAI Scratchpad 0.171 0.097 0.134 [0.129, 0.139] <0.001 31%
Claude-3-5-Sonnet-20240620 Anthropic News with freeze values Scratchpad 0.187 0.083 0.135 [0.129, 0.141] <0.001 32%
Always 0.5 ForecastBench 0.250 0.173 0.212 [0.209, 0.215] <0.001 25%
@@
1. Shows performance on the 1,000 (500 standard, 500 combination) questions in the LLM question set at the 7-, 30-, and 90-day forecast horizons.
2. The full leaderboard is available at www.forecastbench.org. Online results are updated nightly, so may be slightly different than the static version presented here.
3. For resolved questions, predictions are compared against ground truth while for unresolved questions, they are compared to community aggregates.
@@
5. Pct. more accurate than No. 1: The percent of questions where this forecaster had a better overall score than the best forecaster (with rank 1).
6. Pairwise p-value comparing to No. 1 (bootstrapped): The p-value calculated by bootstrapping the differences in overall score between each model and the best forecaster (the group with rank 1) under the null hypothesis that there’s no difference.
@@
Figure 1: Two graphs demonstrating the linear relationship between the Brier scores from Table 2 and (a) Chatbot Arena scores and (b) estimates of training compute. The dotted blue line represents the Superforecasters’ overall Brier score. A red dot with a bootstrapped 95% confidence interval is placed at the intersection of this dotted blue line with the dashed linear fit line to demonstrate the potential intersection of LLM Arena score/training compute and Superforecaster-level forecasting performance. For (b), where available, estimates are from Epoch AI (2024); Llama 3-8b was estimated following https://epoch.ai/blog/estimating-training-compute and we assumed minimum training compute estimates for Claude 2.1, Qwen 1.5 110B and GPT 3.5 Turbo equivalent to Epoch’s estimates for Claude 2, Qwen 1.5 72B and GPT 3.5, respectively. The trend-line in (a) is $y=0.506-0.000298x$ ($R^{2}=0.47$) and in (b) it is $y=0.936-0.014x$ ($R^{2}=0.47$).
@@
1(a) demonstrates the linear relationship between Chatbot Arena scores (see Chiang et al. (2024)) and the overall Brier score from Table 2. We observe a significant correlation ($r=-0.69$, $p=0.002$), indicating that models with higher Arena scores tend to produce more accurate forecasts. The linear relationship implies that LLMs will match superforecaster performance when the Arena score approaches $1398$ (bootstrapped 95% CI: 13351335133513351581158115811581).
1(b) shows the log-linear relationship between estimated training compute and the overall Brier score from Table 2. Projecting out the log-linear relationship, we find that LLMs could match superforecaster performance when training compute approaches $4.43\times 10^{26}$, though there is a large confidence interval (bootstrapped 95% CI: $8.31\times 10^{25}$$1.93\times 10^{28}$) given the marginally significant relationship ($r=-0.69$, $p=0.04$).

### v5

@@ 3.4 Leaderboards
We generate leaderboards that rank models and humans by average overall score. The main leaderboard highlights the top forecasting submission across all questions and can by sorted by performance on the question type (market or dataset) and by resolution status (resolved or unresolved). The leaderboard is updated nightly after scoring forecasts against the latest data, market resolution values, and crowd forecasts.
@@ 3.5 Datasets
Every time we run the public survey described in Section 4, we provide multiple independent forecasts and rationales for every one of the 200 forecast questions and report the accuracy of the median public forecast. See Section B.2.1 for details.
@@
In addition to forecasts and rationales, the superforecasters provide pertinent information about their forecasting process, like search terms used and useful URLs consulted. See Section B.2.2.
@@
Similar to the general public dataset, we ask LLMs to produce forecasts on each of 1,000 forecast questions in the LLM question set. Their rationales are also included in the dataset whenever provided. See Section B.2.3 for details.
@@
In creating the benchmark, we have automated question creation and resolution from all of the sources outlined in Table 1. We provide these as a dataset that can be combined with the forecast datasets mentioned above. See Section B.1 for details.
@@ 4 Human forecaster baseline
To compare LLM forecasting performance to human performance, we ran surveys of two different groups: the general public and superforecasters. We scored each group’s median forecast, treating it as representative of its overall performance.
4.1 General public
We recruited 500 human forecasters via Prolific and advertisements on Facebook to participate as representatives of the general public. These human subjects completed a brief introductory survey to gather demographic information [^8] See Appendix L for an overview of public participant demographics. and evaluate performance on a few forecasting and comprehension tasks. They then completed a one-hour survey containing 20 random questions from the 200-item human question set described in Section 3.2, providing their forecasts and rationales for each question.
The number of responses per question varied to ensure representativeness across categories and sources; at least 40 responses were gathered per question, averaging 49 responses per question.
@@ 4.2 Superforecasters
We recruited 39 superforecasters, who have a strong track record of accurate performance on a diverse set of geopolitical questions, to participate in a 9-day forecasting experiment in which participants were prompted to give their individual forecasts for 20 random questions from the same 200-item human question set described above. Roughly halfway through the 9-day experiment, participants were moved into a group forecasting stage in which we allowed them to see one another’s forecasts and rationales and to communicate about each question. They were also given the opportunity to forecast on questions beyond the 20 questions assigned to them individually.
Because of the lower number of superforecasters, questions generally had fewer responses than in the public survey; a minimum of 3 forecasts were recorded for each question, with an average of 8 responses per question. [^9] See Figure 2 for an example question from the human surveys.
@@ 5 LLM baseline
In this section, we evaluate the forecasting capabilities of LLMs and report on the methodology and results.
@@ 5.1 Methodology
We evaluate a suite of instruction-following chat models without any additional fine-tuning (see Section 2 for details on the models). For each baseline outlined below, we prompt the model to generate a probabilistic forecast that the question will resolve to “Yes.”
@@
    Our first baseline prompts the model zero-shot to generate a forecast directly without generating other content, such as intermediate thinking (Figure 4). By prompting the model to output its forecast directly, we assess raw forecasting capability without sensitivity to prompting strategies.
@@
    Since LLMs’ knowledge is not continuously updated, it is important to provide them with up-to-date information relevant to the question (Zou et al., 2022). Our fourth baseline, scratchpad with news, uses the same scratchpad prompt as above, supplemented with retrieved news articles. The retrieval system is the same as described in Halawi et al. (2024): an LLM generates search queries for a news API, filters articles for relevancy, and summarizes the articles.
@@
    The question sets we provide to LLMs contain what we term freeze values. For market questions these are just the crowd forecast on the market the day the question set was created, as described in Section 3.2. For dataset questions, these are baseline values relevant to the forecasting task. [^10] For example, in a question generated from a Wikipedia page about whether a chess player’s Elo rating will increase by a given date, the freeze value is the chess player’s Elo rating on the question set generation date. An explanation of what the freeze value represents is also provided. Our third baseline is the zero-shot with freeze values. This is simply the zero-shot prompt from Baseline 1 supplemented with the freeze value and an explanation of the freeze value. For examples of the freeze value and its explanation, see Table 8 and Table 9.
@@
    In our final baseline, we aggregate the predictions generated by LLMs into an LLM “ensemble” forecast. We do this as Metaculus (2023) shows that an ensemble of all forecasters consistently outperforms using just the top 5, 10, …, 30 best forecasters (based on past scores).
    To produce the LLM ensemble forecast, we use 3 models (GPT-4o, Claude-3.5-Sonnet, and Gemini-1.5-Pro) and 3 prompts crafted by superforecasters (Figure 6, Figure 7, and Figure 8). This results in 9 forecasts per question. We generate 3 LLM ensemble baselines using the median, geometric mean, and geometric mean of log odds (Satopää et al., 2014). For details, see Appendix E.
@@ 5.2 Results
In Table 2, we show that superforecasters achieve an overall mean Brier score of $0.096$, significantly outperforming both the general public (Brier $=0.121$, $p<0.001$) and the top LLM performer on the 200-item subset (Claude 3.5 Sonnet: Brier $=0.122$, $p<0.001$). [^11] See statistical note in Appendix G. The top-performing LLMs all had access to the crowd forecast on market questions (the “freeze values” from Baselines 4, 5, and 6 above). The top-performing model without access to the crowd forecast on market questions was less accurate than models with access to the human forecast with a Brier score of $0.136$. The comparison between humans and LLMs relies on the 200 questions forecasted by humans, which is a random sub-sample of the 1,000 questions in the question set provided to LLMs (excluding combination questions). [^12] Accuracy measures are based on more than 200 forecasts because human and LLM forecasters submitted multiple forecasts on each dataset question, one for each time horizon. The results presented here include forecasts over the 7-, 30-, 90-, and 180-day time horizons.
@@
Superforecaster median forecast ForecastBench 0.118 0.074 0.096 [0.076, 0.116] 0%
Public median forecast ForecastBench 0.153 0.089 0.121 [0.101, 0.141] <0.001 22%
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Scratchpad 0.138 0.107 0.122 [0.099, 0.146] <0.001 31%
Claude-3-5-Sonnet-20240620 Anthropic News with freeze values Scratchpad 0.142 0.112 0.127 [0.104, 0.150] <0.001 29%
GPT-4-Turbo-2024-04-09 OpenAI Freeze values Zero shot 0.162 0.095 0.128 [0.105, 0.151] <0.001 32%
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Zero shot 0.145 0.117 0.131 [0.103, 0.159] <0.001 31%
GPT-4 OpenAI Freeze values Zero shot 0.167 0.096 0.132 [0.109, 0.155] <0.001 31%
GPT-4o OpenAI News with freeze values Scratchpad 0.162 0.105 0.133 [0.113, 0.154] <0.001 25%
Claude-3-5-Sonnet-20240620 Anthropic Scratchpad 0.138 0.133 0.136 [0.113, 0.158] <0.001 28%
GPT-4o OpenAI Freeze values Scratchpad 0.161 0.113 0.137 [0.115, 0.158] <0.001 27%
@@
1. Shows performance on the 200 standard questions provided in the human question set at the 7-, 30-, 90-, and 180-day forecast horizons. See Table 18 for top 50.
2. The full leaderboard is available at www.forecastbench.org. Online results are updated nightly, so may be slightly different than the version presented here.
3. For resolved market questions, forecasts are compared against ground truth while for unresolved market questions, they are compared to community aggregates.
@@
5. Pairwise $p$-value comparing to No. 1 (bootstrapped): The $p$-value calculated by bootstrapping the differences in overall score between each model and the best
forecaster under the null hypothesis that there’s no difference.
6. Pct. more accurate than No. 1: The percent of questions where this forecaster had a better overall score than the best forecaster.
As a particular failure mode, we find LLMs are significantly worse at combination questions. Although our human surveys did not explicitly ask for forecasts on combination questions, we bound human performance by assuming independence of the component of each combination question. This underestimates human accuracy because a human forecaster predicting the outcome of a combination question could account for dependence between the permuted events. In Table 20, we present this comparison of human and LLM forecasts. We see that LLMs perform poorly on these combination questions, and including them in the benchmark widens the gap between human and LLM performance: superforecasters (Brier $=0.076$) outperform the general public (Brier $=0.096$) and the top LLM (GPT-4o, Brier $=0.130$) significantly. To benchmark the size of this gap in performance, the $0.054$ Brier score gap in performance between superforecasters and GPT-4o is significantly larger than the $0.026$ gap in performance between GPT-4o and GPT-4.
@@
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Scratchpad 0.169 0.078 0.123 [0.117, 0.129] 0%
GPT-4-Turbo-2024-04-09 OpenAI Freeze values Scratchpad 0.172 0.080 0.126 [0.120, 0.132] 0.096 43%
GPT-4o OpenAI Freeze values Scratchpad 0.186 0.069 0.128 [0.122, 0.133] <0.01 43%
Gemini-1.5-Pro Google Freeze values Scratchpad 0.162 0.106 0.134 [0.128, 0.139] <0.001 35%
GPT-4o OpenAI News with freeze values Scratchpad 0.190 0.084 0.137 [0.131, 0.143] <0.001 39%
Gemini-1.5-Pro Google News with freeze values Scratchpad 0.166 0.111 0.139 [0.133, 0.144] <0.001 34%
Claude-3-Opus-20240229 Anthropic Freeze values Zero shot 0.186 0.093 0.139 [0.133, 0.146] <0.001 41%
Qwen1.5-110B-Chat Qwen Freeze values Scratchpad 0.176 0.108 0.142 [0.136, 0.148] <0.001 30%
Claude-3-5-Sonnet-20240620 Anthropic News with freeze values Scratchpad 0.184 0.101 0.143 [0.137, 0.149] <0.001 32%
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Zero shot 0.192 0.094 0.143 [0.136, 0.150] <0.001 42%
@@
1. Shows performance on the 1,000 (500 standard, 500 combination) questions in the LLM question set at the 7-, 30-, 90-, and 180-day forecast horizons. See Table 19
for top 50.
2. The full leaderboard is available at www.forecastbench.org. Online results are updated nightly, so may be slightly different than the version presented here.
3. For resolved market questions, forecasts are compared against ground truth while for unresolved market questions, they are compared to community aggregates.
@@
5. Pairwise $p$-value comparing to No. 1 (bootstrapped): The $p$-value calculated by bootstrapping the differences in overall score between each model and the best
forecaster under the null hypothesis that there’s no difference.
6. Pct. more accurate than No. 1: The percent of questions where this forecaster had a better overall score than the best forecaster.
@@
Figure 1: The graphs show the linear relationship between the Brier scores from Table 2 and (a) Chatbot Arena scores and (b) estimates of training compute. The dotted blue line represents the Superforecasters’ overall Brier score. A red dot with a bootstrapped 95% confidence interval is placed at the intersection of this dotted blue line with the dashed linear fit line to demonstrate the potential intersection of LLM Arena score/training compute and Superforecaster-level forecasting performance. For (b), if estimates from Epoch AI (2024) were not available, we produced estimates following https://epoch.ai/blog/estimating-training-compute. The trend-line in (a) is $y=0.506-0.000298x$ ($R^{2}=0.47$) and in (b) it is $y=0.844-0.01213x$ ($R^{2}=0.41$).
@@
1(a) demonstrates the seemingly linear relationship between Chatbot Arena scores (Chiang et al., 2024) and the overall Brier score from Table 2. We observe a significant correlation ($r=-0.68$, $p=0.003$), indicating that models with higher Arena scores tend to produce more accurate forecasts. The linear relationship implies that LLMs could match superforecaster performance when the Arena score approaches $1406$ (bootstrapped 95% CI: 13461346134613461633163316331633).
1(b) shows the log-linear relationship between estimated training compute and the overall Brier score from Table 2. Projecting out the log-linear relationship, we find that LLMs could match superforecaster performance when training compute approaches $6.49\times 10^{26}$, though there is a large confidence interval (bootstrapped 95% CI: $9.69\times 10^{25}$$8.65\times 10^{28}$) given the marginally significant relationship ($r=-0.67$, $p=0.046$).